OpenStudy (anonymous):

Use the given root to find the solution set of the polynomial x^3+3x^2-8x+10=0,1+i Answer {1 + i, 1 - i, -5} {1 + i, 1 - i, 5} {1 + i, 1 - i, 5i} {-5, 5}

OpenStudy (anonymous):

other root must be the conjugate \(1-i\)

OpenStudy (anonymous):

and the quadratic that gives those two roots is \[x^2-2x+2\] so you can factor that out and find the other root

OpenStudy (anonymous):

huh....

OpenStudy (anonymous):

hmmm ok lets go with the first thing i wrote

OpenStudy (anonymous):

you are given one root at \(1+i\) do you know what the other one must be? and why?

OpenStudy (anonymous):

no and no

OpenStudy (anonymous):

ok the roots of a polynomial with real coefficients come in conjugate pairs, meaning if \(a+bi\) is one root, then so is \(a-bi\) so if \(1+i\) is one root, then so is \(1-i\)

OpenStudy (anonymous):

the reason is that they come like that is from the quadratic formula which has \(x=\frac{-b+\sqrt{b^2-4ac}}{2a}\) and its conjugate \(\frac{-b-\sqrt{b^2-4ac}}{2a}\) you get the complex number and its conjugate if \(b^2-4ac<0\)

OpenStudy (anonymous):

that is how you know that the other root is \(1-i\) now as for the last root, if you did not understand what i wrote above, we already know that answer C and D are not right, there is only one more root, because the polynomial has degree 3 and we have two of them. also it has to be a real number, because the complex ones come in pairs

OpenStudy (anonymous):

so it is either 5 or -5 and i guess you can check which one gives you zero